please in part b graph it with identifying everything. 2. Consumer Theory. Ahns utility function for goods X (pizzas) and Y (cola) is represented as U(X, Y) = 2ln(X)+ln(Y). The pr
Answer #1

U(XY) = 210(X) +inly) Px= $1 Py= $1 m= $12 Consumea will optimise its utility co marimize I his level de utility ata point whUlassinate Date Page but x în Budget line El Pre-x+ ly.y=m x + y = 12 Gyt y = 12 - 3y =12 lys y LIE Suppose there is an Increr- Price Effect = Substitution Effect — Income Effect Substitution Effect ! It refers to change in Relative Price pato keepinPage As shown in the greeph 1 As pecice of x increases, due to substitution Il effect consumer will substitute As a result, n